OpenStudy (anonymous):

Prove that 6 divides ((n^3)-n) whenever n is a non negative integer by using inductions... basis step and inductive step.. Can some one help me please

OpenStudy (anonymous):

we need to prove that 6 divides n(n-1)(n+1)

OpenStudy (anonymous):

how did you break it into that? and from their how do you go about making the basis step and then making the inductive step

OpenStudy (anonymous):

n^3-n=n(n^2-1)=n(n-1)(n+1)

OpenStudy (anonymous):

Okay thank you